What is the kinetic energy in MeV of a ?
Answer #1

tau = gamma tau0

gamma = 1.8E-16/0.84E-16= 2.14


KE = (gamma - 1) m c^2 = (2.14-1)*135=154 MeV
Answer #2

90.0 MeV